OpenStudy (anonymous):

Compute the unit vectors in the direction of |v1| and |v2|, where v1=(2,-6) v2=(-4,7) |v1|= 6.32 |v2|=8.06

jimthompson5910 (jim_thompson5910):

unit vector for v1 u1 = (v1)/(|v1|) unit vector for v2 u2 = (v2)/(|v2|)

jimthompson5910 (jim_thompson5910):

it's equivalent to saying \[\Large u_1 = \frac{1}{|v_1|}*v_1\] \[\Large u_2 = \frac{1}{|v_2|}*v_2\]

OpenStudy (anonymous):

So (2,-6)/6.32 and (-4,7)/8.06?

OpenStudy (anonymous):

How would one use the point/vector over a real number?

jimthompson5910 (jim_thompson5910):

well the portion 1/(|v1|) is a scalar

jimthompson5910 (jim_thompson5910):

so you have a scalar times a vector

jimthompson5910 (jim_thompson5910):

c*<a,b> = <a*c, c*b>

OpenStudy (anonymous):

Ah, so v1 would be (2/6.32, -6/6.32)?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

u1 actually

OpenStudy (anonymous):

My thanks, good sir

jimthompson5910 (jim_thompson5910):

sure thing